OpenStudy (anonymous):

I am using the ratio test to determine convergence or convergence of the series defined by ∂ =10; ∂(sub)n+1 = (5/n)∂n Can someone help?

OpenStudy (anonymous):

\[ \delta_2= \frac{ 10(5)}1\\ \delta_3= \frac{ 10(5^2)}{1*2}\\ \delta_4= \frac{ 10(5^3)}{1*2*3}\\ \] Can you write \[ \delta_n \]

OpenStudy (anonymous):

\[ \delta_n= \frac {10(5^{n-1})} {(n-1)!} \]

OpenStudy (anonymous):

wow this stuff looks complicated what kind of math is this

OpenStudy (anonymous):

\[ \frac {\delta_{n+1}}{\delta_n}=\frac{10\ 5^n}{\frac{\left(10\ 5^{n-1}\right) n!}{(n-1)!}}=\frac{5 (n-1)!}{n!}=\frac{5}{n}-> 0 < 1 \] When n goes to Infinity. Series is convergent.

OpenStudy (anonymous):

It is Calc II, Calc III stuff